OpenStudy (anonymous):

If you decrease the distance between two identical point charges so that the new distance is five time the original distance apart, what happens to the force between them?

OpenStudy (anonymous):

A. it is multiplied by 5 B. it is multiplied by 25 C. it is divided by 5 D. it is divided by 25

OpenStudy (anonymous):

What is the equation for the force and how is it related to the distance?

OpenStudy (anonymous):

i forget :(

OpenStudy (michele_laino):

before decreasing the distance we can write: \[\Large F = K\frac{{{Q_1}{Q_2}}}{{{d^2}}}\] whereas after, we can write: \[\Large {F_1} = K\frac{{{Q_1}{Q_2}}}{{{{\left( {5d} \right)}^2}}}\]

OpenStudy (anonymous):

ok! and so what do i plug in? :/

OpenStudy (michele_laino):

here is next step: \[{F_1} = K\frac{{{Q_1}{Q_2}}}{{{{\left( {5d} \right)}^2}}} = K\frac{{{Q_1}{Q_2}}}{{25{d^2}}} = \frac{1}{{25}}K\frac{{{Q_1}{Q_2}}}{{{d^2}}} = \frac{1}{{25}}F\]

OpenStudy (michele_laino):

\[\Large {F_1} = K\frac{{{Q_1}{Q_2}}}{{{{\left( {5d} \right)}^2}}} = K\frac{{{Q_1}{Q_2}}}{{25{d^2}}} = \frac{1}{{25}}K\frac{{{Q_1}{Q_2}}}{{{d^2}}} = \frac{1}{{25}}F\]

OpenStudy (michele_laino):

F is the initial force, whereas F_1 is the final force

OpenStudy (anonymous):

i cannot see what you wrote on that last part of the equation :/

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

i see 1/25 and then teh rest disappears :(

OpenStudy (anonymous):

@Michele_Laino ?

OpenStudy (michele_laino):

\[\large {F_1} = K\frac{{{Q_1}{Q_2}}}{{{{\left( {5d} \right)}^2}}} = K\frac{{{Q_1}{Q_2}}}{{25{d^2}}} = \frac{1}{{25}}K\frac{{{Q_1}{Q_2}}}{{{d^2}}} = \frac{1}{{25}}F\]

OpenStudy (michele_laino):

the magnitude of the final force is 1/25 of magnitude of the initial force

OpenStudy (anonymous):

oh so our solution is D? divided by 25?

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (anonymous):

yay! thank you!!

OpenStudy (michele_laino):

thank you!! :)